OpenStudy (anonymous):

Solve the following equations. Show all work including calculator-ready answers. (Round to nearest thousandth) 1. 5in(3x - 2) = 15 2. In x + In 4 = 2 3. -7 + In 2x = 4

zepdrix (zepdrix):

\(\large 5\text{in}\) ? is that suppose to be \(\large \sin\) in the first one?

OpenStudy (anonymous):

No it is the LOG

OpenStudy (anonymous):

Then it's ln not in (as in l for lima)

OpenStudy (anonymous):

1. \[5\times \ln{(3x-2)} = 15\]\[\ln{(3x-2)}=3\]\[3x-2=e^3\]Hopefully you can solve that for x. 2. Use the identity\[\ln(a)+\ln(b)=\ln(a\times b)\] 3. \[\ln{2x}=11\]\[2x=e^{11}\]Again, solve for x.